Suppose that the monthly incomes of all employees in Firm A are normally distributed with a mean of $1784.85 and a standard deviation of $199.23. If a random sample of 121 employees is selected from Firm A, what is the probability that the mean of the employees in the sample will be greater than $1809.572632 (approximately)? 



1
Expert's answer
2021-02-03T00:14:38-0500

"\bar{x}=1784.85 \\

σ = 199.23 \\

n = 121 \\

P(X>1809.572632) = P(\frac{x-μ}{σ/ \sqrt{n}} > \frac{1809.572632 -1784.85}{199.23/ \sqrt{121}}) \\

= P(Z>1.365) \\

= 1 -P(Z<1.365)"

From Z-table

"= 1 -0.9138 \\

= 0.0862"

Answer: 0.0862